# Improving Retrieval for RAG based Question Answering Models on Financial Documents

### v1

@@ 1.2 Retrieval Augmented Generation
Thus comes in the retrieval aspect of retrieval augmented generation (RAG), in order to decide which chunk of text to add to the input query for question-answering tasks. We explore the use case of document-based question answering, specifically for finance-specific documents and queries. Suppose one has several reports such as 10-ks or earning transcripts and would like to chat or ask questions on them. For this use case, we need to somehow effectively inject the knowledge of those financial documents into the large language model. In order to prevent hallucinations, some sort of in-context learning approach would be ideal where we can specify via prompts for the model to only refer to the additional context given with the question. * add some sort of like example prompt template here *
@@
Below is a diagram of how a traditional RAG pipeline works. First, the user asks the model a question, for example for financial-based Q-A, they might ask *example of financy question here*. This question is then converted into embeddings or numerical representations of the text. In order to find the relevant context, we do a similarity search (such as cosine similarity) between the query embedding and the embeddings contained in the knowledge hub. The knowledge hub is created from the set of documents we want to chat with. After processing the unstructured documents in some way, the documents are divided into chunks and each chunk is converted to embeddings in the form of a vector. All these embeddings are stored in some sort of database, either locally or in the cloud and they consist of the knowledge hub. The algorithm retrieves the relevant chunk from this hub, and the chunk that is the most similar to the user query is chained to it and given to the LLM as context.
@@ 2.1 Chunking Techniques
When it comes to financial reports, these documents are generally quite long and contain more complicated structures like tables. Documents of the same nature, such as 10-ks also generally follow a specific format that can be indicated by the headings and subheadings of the documents. Due to the special nature of these financial reports, a version of element-based chunking that takes such facts into account could help to further improve retrieval. In this paper, we will be following strategies utilized in the Unstructured paper  [^6] Antonio Jimeno Yepes, Yao You, Jan Milczek, Sebastian Laverde, and Renyu Liu. Financial Report Chunking for Effective Retrieval Augmented Generation, 2024. arXiv:2402.05131. - if a title element is found, a new chunk is started and if a table element is found, a new chunk is started, preserving the entire table. Such strategies will help ensure that each context will return just enough context to effectively answer the user’s question.
@@ 2.2 Query Expansion
Query Expansion, or query transformation refers to the idea of changing the question fed into the RAG pipeline to find relevant chunks based on more than the user’s original question. The idea behind this is that oftentimes, the user’s question may not explicitly contain all the information to indicate the algorithm where in the document it should look for the context, especially it is via a cosine similarity search. Sometimes, the user’s question alone can mislead the algorithm to look in an incorrect area and therefore return an irrelevant chunk. (An example of that here would be cool). For more complicated questions and documents, more reasoning steps are required to know where to look and basic RAG pipelines lack this completely.
@@ 2.4 Re-ranking Algorithms
Re-ranking algorithms is a method to prioritize the relevance over the similarity of the chunks. Essentially, a method like cosine similarity might rank the top 10 chunks, but a separate algorithm will re-rank the algorithms to be based on relevance, and then the top one or two chunks after the re-ranking will be augmented as context to the input query. Cohere’s re-ranking algorithm is a popular one and it along with others uses additional machine learning and natural language processing techniques to further evaluate relevance beyond a similarity search.  [^12] Cohere. Rerank, 2024. https://cohere.com/rerank
@@ 2.5 Fine-tuning Embedding Algorithms
Embedding algorithms are what convert text into numerical representations, and play a crucial role in RAG pipelines. It is possible to fine-tune embedding algorithms based on domain-specific knowledge to enhance retrieval in that specific domain. Embeddings can also be dynamic where they adapt when words have slightly different meanings based on the context. OpenAI’s embeddings are an example of dynamic embeddings that change based on the context. Domain-specific fine-tuning, however, requires access to datasets that consist of queries, the corpus of text, and relevant documents about that specific domain.  [^13] Yunfan Gao, Yun Xiong, Xinyu Gao, Kangxiang Jia, Jinliu Pan, Yuxi Bi, Yi Dai, Jiawei Sun, Qianyu Guo, Meng Wang, and Haofen Wang. Retrieval-Augmented Generation for Large Language Models: A Survey, 2024. arXiv:2312.10997.
@@ 3.2 Answer Accuracy
4 Dataset
To assess the performance of a model with RAG, we utilized the FinanceBench benchmark for question and answering models developed by Patronus AI. This benchmark consists of 10,231 questions about publicly traded companies in the USA, covering a range of financial documents such as 10Ks, 10Qs, 8Ks, and Earnings Reports released between 2015 and 2023.
@@ 5 Impact
6 Results
Due to the dynamic nature of the evolving AI landscape, we are not publicly disclosing the results of the techniques at this time. Feel free to contact our team, as we can provide results on a case by case basis.

### v2

@@ 1.2 Retrieval Augmented Generation
Thus comes in the retrieval aspect of retrieval augmented generation (RAG), in order to decide which chunk of text to add to the input query for question-answering tasks. We explore the use case of document-based question answering, specifically for finance-specific documents and queries. Suppose one has several reports such as 10-ks or earning transcripts and would like to chat or ask questions on them. For this use case, we need to somehow effectively inject the knowledge of those financial documents into the large language model. In order to prevent hallucinations, some sort of in-context learning approach would be ideal where we can specify via prompts for the model to only refer to the additional context given with the question. For example by the following as a systems prompt: ”””You are a financial chatbot trained to answer questions based on the information provided in 10-K documents. Your responses should be directly sourced from the content of these documents.””” and this as a user prompt: f”Context: context Query: query Answer:”
@@
Below is a diagram of how a traditional RAG pipeline works. First, the user asks the model a question, for example for financial-based Q-A, they might ask ”What is the FY2022 capital expenditure amount (in USD millions) for Company X?” This question is then converted into embeddinga or numerical representations of the text. In order to find the relevant context, we do a similarity search (such as cosine similarity) between the query embeddings and the embeddings contained in the knowledge hub. The knowledge hub is created from the set of documents we want to chat with. In this case, the documents might be a set of financial documents about various companies. After processing the unstructured documents in some way, the documents are divided into chunks and each chunk is converted to embeddings in the form of a vector. All these embeddings are stored in some sort of database, either locally or in the cloud and they consist of the knowledge hub. The algorithm retrieves the relevant chunk from this hub, and the chunk that is the most similar to the user query is chained to it and given to the LLM as context.
Figure 1: Retreival Augmented Generation Architecture
@@ 2.1 Chunking Techniques
When it comes to financial reports, these documents are generally quite long and contain more complicated structures like tables. Documents of the same nature, such as 10-ks also generally follow a specific format that can be indicated by the headings and subheadings of the documents. Due to the special nature of these financial reports, a version of element-based chunking that takes such facts into account could help to further improve retrieval. A paper by Unstructured proposed a novel way of breaking up documents into manageable sizes via element-based chunking.  [^6] Antonio Jimeno Yepes, Yao You, Jan Milczek, Sebastian Laverde, and Renyu Liu. Financial Report Chunking for Effective Retrieval Augmented Generation, 2024. arXiv:2402.05131. - if a title element is found, a new chunk is started and if a table element is found, a new chunk is started, preserving the entire table. Such strategies will help ensure that each context will return just enough context to effectively answer the user’s question.
Other chunking methods include leveraging NLP techniques to implement recursive, character based chunking or semantic chunking. The nature of semantic or recursive chunking preserves the context of the document much better than character based or uniform chunking methods but become more complex to implement. A step about this is using something called agentic chunking where another LLM is prompted to break up the text in the best way possible. While this methods would seem to have the best contextual understanding, it is the most computationally expensive and often runs into the issue of exceeding the token size limit for the embedding models.
@@ 2.2 Query Expansion
Query Expansion, or query transformation refers to the idea of changing the question fed into the RAG pipeline to find relevant chunks based on more than the user’s original question. The idea behind this is that oftentimes, the user’s question may not explicitly contain all the information to indicate the algorithm where in the document it should look for the context, especially it is via a cosine similarity search. Sometimes, the user’s question alone can mislead the algorithm to look in an incorrect area and therefore return an irrelevant chunk. For more complicated questions and documents, more reasoning steps are required to know where to look and basic RAG pipelines lack this completely.
@@ 2.4 Re-ranking Algorithms
Re-ranking algorithms is a method to prioritize the relevance over the similarity of the chunks. Essentially, a method like cosine similarity might rank the top 10 chunks, but a separate algorithm will re-rank the algorithms to be based on relevance, and then the top one or two chunks after the re-ranking will be augmented as context to the input query. Cohere’s re-ranking algorithm is a popular one and it along with others uses additional machine learning and natural language processing techniques to further evaluate relevance beyond a similarity search. In this paper we utilized a cross-encoder model to score how relevant each retreived chunk was to the question and re ranked the chunks based off the scores. The nature of the architecture suggests that cross encoder models have higher accuracy than bi-encoder models for this specific use case.  [^12] Cohere. Rerank, 2024. https://cohere.com/rerank
@@ 2.5 Fine-tuning Embedding Algorithms
Embedding algorithms are what convert text into numerical representations, and play a crucial role in RAG pipelines. It is possible to fine-tune embedding algorithms based on domain-specific knowledge to enhance retrieval in that specific domain. Embeddings can also be dynamic where they adapt when words have slightly different meanings based on the context. OpenAI’s embeddings are an example of dynamic embeddings that change based on the context. Domain-specific fine-tuning, however, requires access to datasets that consist of queries, the corpus of text, and relevant documents about that specific domain. In this paper, we are focusing on zero-shot methods to improve retrieval that require no additional training data. However future work will investigate how additional annotations/data can possible improve RAG via fine tuning embedding algorithms.  [^13] Yunfan Gao, Yun Xiong, Xinyu Gao, Kangxiang Jia, Jinliu Pan, Yuxi Bi, Yi Dai, Jiawei Sun, Qianyu Guo, Meng Wang, and Haofen Wang. Retrieval-Augmented Generation for Large Language Models: A Survey, 2024. arXiv:2312.10997.
@@ 3.2 Answer Accuracy
4 Results
To assess the performance of a model with RAG, we utilized the FinanceBench benchmark for question and answering models developed by Patronus AI. This benchmark consists of 10,231 questions about publicly traded companies in the USA, covering a range of financial documents such as 10Ks, 10Qs, 8Ks, and Earnings Reports released between 2015 and 2023. However only 150 rows were publicly avaiable, so that is what was used in this study
@@ 4 Results
The FinanceBench dataset was tested for the general RAG model called base RAG. In the base case, the RAG model was not given additional context, meaning the model had to retrieve context information on its own. On the contrary, another experiment was tested, we will call this Fake RAG. Fake RAG uses the same LLM model but is given the correct context to answer the original question, specifically using context information from FinanceBench. We also have results for the whole dataset for the query expansion techniques and a reranking approach using a cross encoder. There are some initial results for testing different chunking techniques, however because of compute limitations as well as it exceeding the token limit of the model, we could not get a comprehensive metric of the whole dataset and therefore cannot compare it directly with the other methods. The preliminary results for those chunking methods will be available in the github repository for anyone interesting in seeing the initial results and code. While we had started to experiment with some metadata tagging techniques, there was not a clear way to integrate such with the existing pipeline so results for that technique are not currently available, but the corresponding repository will be updated when we have run successful experiments.
As you can see from the plot and table, the ”fake rag” approach performed the best with an average LLM evaluation score of 0.573. This was expected to have the best results as the LLM was given the context that was required to answer the question correctly. However, as you can see from the table, this method had the highest standard deviation or variation in the score. When looking through the data, it seems that when given the right context, the LLM answered the question correctly, or got it completely wrong. So while giving the correct context significantly improves the accuracy of the model, there are still limitation swith the LLM itself, even when using the state of the art generator like GPT-4o, when answering highly complex and domain specific questions.
The method with the lowest accuracy was the base rag case, with an average LLM evaluation score of 0.204. The same LLM and prompt template was used for each method, only the way the context was retrieved differed. This means that the difference in the context fed into the model contributed to a gap in accuracy of over 35 percent, underscoring the importance of retrieving correct texts, especially for longer more complex documents. The zero shot methods of query expansion and re-ranking seem to have some improvement on the accuracy with LLM evaluation scores of 0.24 and 0.256 respectively. While this is no where near the accuracy of the fake rag results when it is fed the correct context, its improvement over the base rag case is not insignificant. From this, we can conclude that zero shot methods to enhance retrieval might be in the step in the right direction to improve accuracy, but not nearly enough to make robust systems with very high accuracy. In order to to achieve that goal, other techniques, perhaps those that require some additional training or fine tuning of some sorts can be utilized.
Figure 3: Aggregate Results Plot
Figure 4: Aggregate Results Table